OpenStudy (anonymous):

Just an answer please! One leg of a right triangle is 12, and the hypotenuse is 20. The length of the other leg is___

OpenStudy (anonymous):

@amistre64

OpenStudy (anonymous):

C^2=a^2+b^2 . The pythagorean theorem. You know C. You know a. Find b.

OpenStudy (anonymous):

We are not here to provide answers. You will never learn by copying down the answers without understanding what you are doing.

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

20^2 + 12^2 = Answer?

OpenStudy (anonymous):

so 544 is the answer? @Dido525

OpenStudy (anonymous):

No. In the Pythagorean theorem C is the hypotenuse and a, b are the legs of the triangle.

OpenStudy (anonymous):

\[a^2 + b^2 = c^2\] с is the hypotenuse. Make sure that you plug in the numbers properly

OpenStudy (anonymous):

So its 23 ?

OpenStudy (anonymous):

No. Show us your work.

OpenStudy (anonymous):

Okay 20^2 + 12^2 = 544 Squared root of 544 is 23.32

OpenStudy (anonymous):

Careful! You plugged in the numbers wrong.

OpenStudy (anonymous):

What? How?

OpenStudy (anonymous):

The Hypotenuse is the c value. You accidiently said the hypotenuse is a.

OpenStudy (anonymous):

The formula is c^2=a^2+b^2 . C is the hypotenuse. a is the leg of a triangle B is the other length

OpenStudy (anonymous):

Leg sorry not length.

OpenStudy (anonymous):

on my lesson it says a^2+b^2=c^2

OpenStudy (anonymous):

Its the same thing written backwards :p .

OpenStudy (anonymous):

I just chaged around where the terms are.

OpenStudy (anonymous):

But it dosent matter. A^2+b^2=c^2

OpenStudy (anonymous):

You know the c value which is the hypotenuse. You know the a value. You are trying to find b.

OpenStudy (anonymous):

Do you understand?

OpenStudy (anonymous):

then what the heck, you said C^2= blah blah right? So how did i plug it in wrong? 20^2 + 12^2 = 544

OpenStudy (amistre64):

a=12, c=20, find b

OpenStudy (amistre64):

12^2 + b^2 = 20^2

OpenStudy (anonymous):

I didnt want to give it away :/ .

OpenStudy (raden):

b^2 = 20^2 - 12^2 b^2 = (20+12)(20-12) solve for b

OpenStudy (anonymous):

$$ a = \sqrt{20^2 - 12^2} = \sqrt{400 - 144} = \sqrt{256}= 16 $$

OpenStudy (anonymous):

THANK YOU! :D

OpenStudy (anonymous):

Dont give her the answer Sigh...

OpenStudy (anonymous):

@mikelo40 it's better to solve problem yourself then to just copy answers from here or book, take your notes from the class, consult your teacher for problems with understanding, math is realy fun when you get it! good luck :)

OpenStudy (anonymous):

Thanks for that at least.